# DNABERT-2: Efficient Foundation Model and Benchmark For Multi-Species Genome

### 26 Jun 2023 v1

@@ A.1 All Experiment Results
A.2 Hyperparameters
This section presents the hyperparameters we used in the fine-tuning stage on each model. Table 5 shows the number of training steps we used for each task. We use AdamW [Loshchilov and Hutter, 2019] as optimizer. We keep most of the other hyperparameters the same for all the models across all the datasets, including a batch size of $32$, a warmup step of $50$, and a weight decay of $0.01$. For DNABERT and DNABERT-2, we perform standard fine-tuning with a learning rate of $3e\texttt{-}5$, while for the Nucleotide Transformers, we perform parameter efficient fine-tuning (PEFT) using Low-Rank Adaptation (LoRA) with a learning rate of $1e\texttt{-}4$, a LoRA alpha of $16$, a LoRA dropout of $0.05$, and a LoRA $r$ of $8$. The hyperparameters are selected based on grid searches over commonly used ones in preliminary experiments.
EMP TF-M CVC TF-H PD-tata PD-other CPD-tata CPD-other SSP
Num. Epochs 3 1k steps 8 3 10 4 10 4 5
Table 5: The number of training steps we used for the following tasks: Epigenetic Marks Prediction (EMP), Transcription Factor Prediction on the Human genome and the Mouse genome (TF-H and TF-M), Covid Variants Classification (CVC), tata dataset of Promoter Detection (PD-tata), notata and all datasets of Promoter Detection (PD-other), tata dataset of Core Promoter Detection (CPD-tata), notata and all datasets of Core Promoter Detection (CPD-other), and Splice Site Prediction (SSP).
A.3 Preliminary Experiments on Nucleotide Transformer
Since there is no official fine-tuning code of Nucleotide Transformer [Lopez et al., 2023], we use its open-sourced checkpoints in Huggingface Modelhub [^2] https://huggingface.co/InstaDeepAI and train it with our code base using LoRA. For a fair comparison with this model, in this section, we present preliminary experiments that compare the results reported in their paper with the performance of this model under our implementation. We select the epigenetic marks prediction task for benchmarking since it is the only shared task among Lopez et al. [2023] and GUE. The task contains $10$ datasets. For each dataset, we randomly split it into training and test sets with a ratio of 9:1. As shown in Table 6, our LoRA implementation leads to slightly better results than the results reported in the original paper, making our comparison to the model fair and convincing despite the fact that we do not have access to its official fine-tuning implementation.
@@ A.3 Preliminary Experiments on Nucleotide Transformer
Table 6: This table presents the performance of the Nucleotide Transformer on ten datasets of epigenetic marks prediction on the Yeast genome. As shown in the table, our implementation achieves better performance than the results reported in the paper, indicating the fairness of comparison in our experiments. *: Results taken from Lopez et al. [2023].

### 18 Mar 2024 v2

@@ A.1 All Experiment Results
Table 6 presents the evaluation results of DNABERT-2 and baselines on the GUE benchmark.
A.2 Implementation
We pre-train DNABERT-2 with the Masked Language Modeling (MLM) loss with a mask ratio of $15\%$. Notably, we independently mask every token instead of masking spans of continuous tokens like Ji et al. (2021). We use a batch size of $4096$ and a max sequence length of $128$. We train the model for $500000$ steps using the AdamW (Loshchilov & Hutter, 2019) optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.98$, $\epsilon=1e\texttt{-}6$ and weight decay of $1e\texttt{-}5$. The learning rate linearly increases from 0 to $5e\texttt{-}4$ during the first $30000$ steps while linearly decreasing to 0 in the last $470000$ steps.
A.3 Hyperparameters
This section presents the hyperparameters we used in the fine-tuning stage on each model. Table 7 shows the number of training steps we used for each task. We use AdamW (Loshchilov & Hutter, 2019) as optimizer. We keep most of the other hyperparameters the same for all the models across all the datasets, including a batch size of $32$, a warmup step of $50$, and a weight decay of $0.01$. For DNABERT and DNABERT-2, we perform standard fine-tuning with a learning rate of $3e\texttt{-}5$, while for the Nucleotide Transformers, we perform parameter efficient fine-tuning (PEFT) using Low-Rank Adaptation (LoRA) with a learning rate of $1e\texttt{-}4$, a LoRA alpha of $16$, a LoRA dropout of $0.05$, and a LoRA $r$ of $8$. The hyperparameters are selected based on grid searches over commonly used ones in preliminary experiments. The pre-training stage takes approximately 14 days using eight Nvidia RTX 2080Ti GPUs. To train the model, we used the Transformers library (Wolf et al., 2020) and the Composer library (Team, 2021).
EMP TF-M CVC TF-H PD-tata PD-o CPD-tata CPD-o SSP
Epochs 3 1k 8 3 10 4 10 4 5
Table 7: The number of training steps we used for the following tasks: Epigenetic Marks Prediction (EMP), Transcription Factor Prediction on the Human genome and the Mouse genome (TF-H and TF-M), Covid Variants Classification (CVC), tata dataset of Promoter Detection (PD-tata), notata and all datasets of Promoter Detection (PD-o), tata dataset of Core Promoter Detection (CPD-tata), notata and all datasets of Core Promoter Detection (CPD-o), and Splice Site Prediction (SSP). In the task of Transcription Factor Prediction on the Mouse genome, we train the model for 1000 steps on each dataset.
A.4 Preliminary Experiments on Nucleotide Transformer
Since there is no official fine-tuning code of Nucleotide Transformer (Dalla-Torre et al., 2023), we use its open-sourced checkpoints in Huggingface Modelhub [^2] https://huggingface.co/InstaDeepAI and train it with our code base using LoRA. For a fair comparison with this model, in this section, we present preliminary experiments that compare the results reported in their paper with the performance of this model under our implementation. We select the epigenetic marks prediction task for benchmarking since it is the only shared task among Dalla-Torre et al. (2023) and GUE. The task contains $10$ datasets. For each dataset, we randomly split it into training and test sets with a ratio of 9:1. As shown in Table 8, our LoRA implementation leads to slightly better results than the results reported in the original paper, making our comparison to the model fair and convincing despite the fact that we do not have access to its official fine-tuning implementation.
@@ A.4 Preliminary Experiments on Nucleotide Transformer
Table 8: This table presents the performance of the Nucleotide Transformer on ten datasets of epigenetic marks prediction on the Yeast genome. As shown in the table, our implementation achieves better performance than the results reported in the paper, indicating the fairness of comparison in our experiments. *: Results taken from Dalla-Torre et al. (2023).
A.5 Comparing DNABERT-2 with other baselines
We also compare DNABERT-2 with other baselines, including HyenaDNA (Nguyen et al., 2024), a convolutional neural network (CNN) designed for DNA classification by Grešová et al. (2023), and DNABERT-2 model without pre-training. We implement HyenaDNA with their official implementation https://huggingface.co/LongSafari/hyenadna-medium-450k-seqlen-hf on HuggingFace and Huggingface Trainer. We implement CNN with official code on GitHub. For both papers, we use the default hyperparameter described in the papers, and we follow our own evaluation setting (e.g., use the same set of hyperparameters across all the datasets and test on the checkpoint with the lowest validation loss).
As shown in Table 9, DNABERT-2 consistently outperforms the baselines. We note that using the official checkpoint of HyenaDNA on Huggingface does not achieve the same level of performance on the EMP task as reported in their paper. This could be due to multiple reasons: 1) the reported results are not based on the open-source checkpoints, 2) the Huggingface trainer does not train the model properly, and 3) more careful hyperparameter searches are needed.
Epigenetic Marks Prediction
H3 H3K14ac H3K36me3 H3K4me1 H3K4me2 H3K4me3
HyenaDNA 67.17 31.98 48.27 35.83 25.81 23.15
CNN 61.52 29.73 38.60 26.06 25.76 20.52
DNABERT-2 w/o PT 57.36 32.25 38.07 23.80 28.57 18.12
DNABERT-2 78.27 52.57 56.88 50.52 31.13 36.27
Epigenetic Marks Prediction Promoter Detection
H3K79me3 H3K9ac H4 H4ac all notata tata
HyenaDNA 54.09 50.84 73.69 38.44 47.38 52.24 5.34
CNN 46.30 40.03 62.34 25.54 75.78 85.14 70.30
DNABERT-2 w/o PT 51.71 44.38 66.73 30.07 76.49 84.07 34.11
DNABERT-2 67.39 55.63 80.71 50.43 86.77 94.27 71.59
Transcription Factor Prediction (Human) Core Promoter Detection
0 1 2 3 4 all notata tata
HyenaDNA 62.3 67.86 46.85 41.78 61.23 36.95 35.38 72.87
CNN 53.95 63.20 45.22 29.84 61.48 58.07 60.09 69.33
DNABERT-2 w/o PT 59.35 59.27 46.13 31.36 60.14 54.58 58.26 47.15
DNABERT-2 84.38 87.16 83.99 79.85 87.80 69.47 68.04 74.17
Transcription Factor Prediction (Mouse) Virus Splice
0 1 2 3 4 Covid Reconstruct
HyenaDNA 35.62 80.50 65.34 54.20 19.17 23.27 72.67
CNN 0 31.14 59.74 63.15 45.48 27.18 22.23 76.79
DNABERT-2 w/o PT 30.55 63.74 59.92 24.12 27.20 69.76 46.80
DNABERT-2 81.60 92.67 92.38 84.91 76.10 71.02 84.99
Table 9: Benchmark DNABERT-2 with HyenaDNA (Nguyen et al., 2024), CNN (Grešová et al., 2023), and DNABERT-2 without pre-training on GUE.
A.6 Ablation Study on Tokenization Method
In this section, we present an ablation study on the tokenization method: the overlapping K-mer tokenization used in DNABERT v.s. the BPE tokenization used in DNABERT-2. We train two DNABERT-2 variants on the same training data and with the same model architecture and hyperparameters. We use a batch size of $4096$, max sequence length of $128$, and train each model for $120,000$ steps. Since the vocab size of BPE is 4096, to keep the number of parameters similar enough, we use 6-mer tokenization, which results in 4101 tokens. We then evaluate the two variants on the GUE benchmark. Table 10 presents the results. As shown in the Table, the variant trained with the BPE tokenizer outperforms the K-mer counterparts on 21 out of 28 datasets, with an average score of 65.33 and 60.92, respectively. As a result, we empirically show that BPE leads to better performance than k-mer after being pre-trained and finetuned with the same data, architecture, and hyperparameters, showing its data efficiency. Also, as shown in Table 6 of the paper, BPE also leads to 3-4 times less computational costs than K-mer.
Epigenetic Marks Prediction
Tokenizer H3 H3K14ac H3K36me3 H3K4me1 H3K4me2 H3K4me3
K-mer 74.62 42.71 47.26 39.66 25.33 27.43
BPE 77.08 55.60 57.25 45.51 40.83 42.57
Epigenetic Marks Prediction Promoter Detection
Tokenizer H3K79me3 H3K9ac H4 H4ac all notata tata
K-mer 61.03 49.35 78.61 37.14 83.78 92.65 57.75
BPE 66.01 56.79 80.07 54.19 85.57 92.55 60.85
Transcription Factor Prediction (Human) Core Promoter Detection
Tokenizer 0 1 2 3 4 all notata tata
K-mer 67.99 67.06 59.45 50.24 72.80 74.91 69.23 74.91
BPE 66.99 70.98 61.40 55.10 71.31 66.28 67.99 72.73
Transcription Factor Prediction (Mouse) Virus Splice
Tokenizer 0 1 2 3 4 Covid Reconstruct
K-mer 48.96 81.69 81.71 63.17 42.83 62.16 77.90
BPE 48.01 81.86 82.98 73.22 46.15 69.75 79.62
Table 10: This table presents the ablation study on tokenization methods.
@@ B.1 Multi-Species Genome for Pre-Training
Invertebrate Brachionus Rubens 1327.37
Ceroptres Masudai 12.95
Cotesia Typhae 1866.62
Croniades Pieria 3889.85
Drosophila Athabasca 1221.16
Emesis Russula 4848.08
Hydra Oligactis 12597.75
Meganola Albula 3604.25
Oscheius 383.21
Rutpela Maculata 20213.33